The current market price for XYZ is $179 per share. Initial margin is 50%, maintenance margin is 35% and margin interest is 1.75% per year. XYZ pays annual cash dividends of $5.95 per share. You believe the stock price will increase over the next year and wish to trade long using margin. At what price would you receive a margin call? Enter your answer rounded to two decimal places. Do not enter $ or comma in the answer box. For example, if your answer is $12,300.456 then enter as 12300.46 in the answer box.
Answer #1

Amount borrowed is calculated below:

Amount\, borrowed=Share\, price\times \left ( 1-Initial\, margin \right )

Amount\, borrowed=179\times \left ( 1-0.50 \right )

Amount\, borrowed=89.50

The price of the investment for margin call is calculated below:

Maintenance\, margin=1-\frac{Amount\, borrowed}{Investment\, value}

0.35=1-\frac{89.50}{Investment\, value}

\frac{89.50}{Investment\, value}=1-0.35

\frac{89.50}{1-0.35}=Investment\, value

137.69=Investment\, value

Individual will receive a margin call at a price of $137.69..